# Block-corrected Modularity for Community Detection

### v1

@@ I Introduction
Here, we take the opposite approach: for a given network data accompanied by node metadata, we discount the effect of known attributes on network structure in order to isolate the effect of unknown attributes. This amounts to asking whether there exists any community structure beyond the known attributes. Factoring out the known attributes has been previously accomplished in spatial networks, for example, by adapting the definition of modularity to discount the effects of spatial embedding [17]. In a mobile phone network where calls are predominantly placed among persons in close physical proximity, nullifying this physical proximity of callers came to reveal linguistically homogeneous communities, communities of callers speaking the same language.
@@ III.1 Simple network model with hidden communities
where edge probability is proportional to in-degree and out-degree of nodes, normalized by the total number of edges. We analytically compute modularity values of specific partitions with respect to both known and unknown structures to assess which partition is preferred by the block-corrected modularity and the directed modularity.
@@
The derived modularity expressions for $Q^{\mathrm{dir}}_{y}$, $Q^{\mathrm{bc}}_{y}$ and $Q^{\mathrm{dir}}_{x}$ have a similar form to the E-I index [41] which equates to the number of internal links of group members subtracted from the number of external links divided by their summation. Similar to the E-I index modularity values $Q_{x}$ and $Q_{y}$ are measures of the dominance of in-group over out-group edges. As per Eq. (4) and Eq. (5) for a given binary attribute the more segregated its two membership groups are, the larger its corresponding modularity value is. The block-corrected modularity evidently rejects the known attribute community structure ($Q^{\mathrm{bc}}_{x}=0$) regardless of how well-separated and embedded its two comprising groups, i.e., the nodes with $x=0$ and those with $x=1$, are. In contrast, the directed null model’s ability to recover the unknown structure depends on the separatedness of the corresponding attribute’s groups. If the separation of the two groups under known attribute $x$ is more pronounced than those of the unknown attribute $y$, the modularity value $Q_{x}$ is larger than $Q_{y}$, in which case a modularity optimization partitioning algorithm will dismiss the community structure driven by the unknown attribute $y$ in favor of partitioning the network based on the known attribute $x$. The details of the derivations are outlined in Appendix. A.
@@ III.2 Communities in temporal networks
A null model is characterized by its edge probability $P_{ij}$ which describes a generative network model for networks maximally similar to an observed network with no community structure present. Null models may nonetheless exhibit structures different to those existing in the observed network which can speak to their efficacy as a generative model. With citation networks in mind, we evaluate the performance of our proposed block null model against two previously proposed null models: the directed null model [40], which we already introduced in the previous subsection, and the DAG null model [14, 35]. The directed and the DAG null models are designed to operate on directed networks with the latter specifically devised with citation networks in mind as they are often represented by directed acyclic graphs [^1] Real-world citation networks, however, are not always acyclic, requiring pre-processing measures to ensure the acyclic nature holds.. While the directed null model preserves only in- and out-degrees of nodes, the DAG null model additionally encodes two parameters $\mu_{t}$ and $\lambda_{t}$ per time layer $t$ describing the edge connectivity between layers:
@@ III.2.2 Temporal connectivity pattern in null models
To assess this, we define a single generalized $(bt_{\mathrm{max}})\times(bt_{\mathrm{max}})$ connectivity matrix $\mathbf{M}$ for each temporal network, where $M_{(bt_{i}+b_{i})(bt_{j}+b_{j})}$ is the fraction of existing edges (out of all possible edges) from community $b_{i}$ at time $t_{i}$ to community $b_{j}$ at time $t_{j}$. We then compare this to the connectivity matrix of the realizations of the same network passed through block-corrected, DAG and directed null models.
@@
Figure 3: (a) Schematic of the skewed generative network model with two planted communities and a temporal structure where only nodes in adjacent blocks are connected. In addition, there are long links from the last block to the first. (b) Averaged connectivity matrices $\mathbf{M}$ of an ensemble of 100 network realizations of the block-corrected, DAG, and directed null models. The connectivity matrix of the original network, generated using the skewed model (Eq. (10)), is labeled “Original”. The skewed model network has $b=2$ planted communities, $t_{\mathrm{max}}=10$ temporal layers, $N=100$ nodes per temporal layer, mean in-group degree $\langle k^{\mathrm{ing}}\rangle=10$ and out-group degree $\langle k^{\mathrm{outg}}\rangle=8$. (c) The distribution of temporal edge lengths for all three network models: one with a skewed distribution of edge lengths (Eq. (10)), one with an exponential distribution (Eq. (11)), and lastly one with a power law edge length distribution (Eq. (12)). The three observed networks for skewed, exponential, and power law in panel (c) has $t_{\mathrm{max}}=50$ temporal blocks, i.e. time layers, with a fixed number of nodes ($N=100$) per temporal block. The number of planted communities is $b=2$ and each node has a mean in-group degree $\langle k^{\mathrm{ing}}\rangle=10$ and out-group degree $\langle k^{\mathrm{outg}}\rangle=8$. Edge length distributions for the drawn networks from the three null models are shown in their respective columns.
Averaged connectivity matrices of the realizations for the block-corrected, DAG, and directed null models are shown in the following three panels. The temporal connectivity patterns produced in our block-corrected model closely resemble that of the original skewed network, whereas the DAG model smooths the temporal connectivity pattern into an exponential decay. The directed model, on the other hand, assumes almost all community-time pairs are equally likely to be connected.
@@
Figure 3(b) shows that both the block-corrected and the DAG models recover the exponential edge length distribution observed in the original network. This is not the case, however, when the observed edge lengths follow a more heavily skewed distribution. The DAG model falls short of reproducing the distribution, producing an exponential edge length decay instead, while the block-corrected generates an ensemble whose edge lengths closely follow the observed distribution. Similarly, under a power law distribution for the edge lengths, the DAG model generates networks with exponential edge length distributions instead, whereas the original power law distributions is recreated by the block-corrected. In all three generative network models with exponential and skewed distributions, the directed model proves ineffective at capturing the nature of the distributions, producing almost constant temporal edge length distributions only affected by the finite number of the temporal blocks $t_{\mathrm{max}}$.
@@ III.2.3 Susceptibility to erroneous partitioning
As we will describe in Sec. V, heavy-tail distribution of edge lengths is observed in certain real-world systems, including in citation network. As we saw in Fig. 3(b), while our block null model can replicate the temporal connectivity pattern induced by a power law edge length distribution, the two other null models do not reproduce the underlying distribution. To understand the implications of this, we directly measure the ability of each modularity definition to recover the planted partitions in the space of all possible bipartitions of the network.
The quality of a partition can be quantified by the adjusted Rand score between that and the planted partition. A score of zero represents performance no better than random labeling of nodes and one is a complete match between the candidate partitioning and the true one. To simplify this comparison, we assume the number of planted communities is known in all cases.
Figure 5: Temporal homogeneity and labeling accuracy for the temporal planted partition model with power law edge length distribution. Adjusted Rand scores between discovered communities and the true partitions (a) and the Shannon entropy of the discovered communities (b) for an ensemble of 50 synthetic networks with two planted communities over a range of exponents of power law edge length distributions (see Eq. (12)). Adjusted Rand score of zero indicates the performance is no better than random chance and higher values of entropy indicate more even distribution of communities across time layers. The horizontal dashed pink line represents the analytical entropy for a uniform distribution $S=\log_{2}(t_{\mathrm{max}})\approx 7.64$ while the horizontal dashed brown line corresponds to the analytical entropy for an equal vertical cut $S=\log_{2}(t_{\mathrm{max}}/2)\approx 6.64$. In panel (c) fraction of nodes in each of the two discovered communities are plotted across time layers for three networks with $\gamma=-1.4$, $\gamma=-1.9$, $\gamma=-2.0$. The synthetic networks are generated with a fixed number of nodes ($N=200$) per time layer ($t_{\mathrm{max}}=200$). Each node has an in-group degree $\langle k^{\mathrm{ing}}\rangle=8$ and out-group degree $\langle k^{\mathrm{outg}}\rangle=4$. Edges probabilities decrease as $P(\Delta t)\propto\Delta t^{\gamma}$. All three models are coded to find exactly two communities. Error bars in panels (a) and (b) show 95% confidence interval.
@@
The DAG modularity, however, alternatively assigns the majority of the nodes of groups of time layers to one of the two communities for smaller values of $\gamma$. This behavior, which is not reminiscent of the planted (original) communities, can already be seen at $\gamma=-2.0$ and seem to aggravate for more heavy-tailed distributions. This, again, corroborates the observations of drop in entropy in Fig. 5(b). As hypothesized before, the directed modularity makes a vertical cut for all three values of $\gamma$ visualized in Fig. 5(c). The cut seems to become more gradual as the magnitude gets smaller, i.e., the heavy tail characteristic of the edge length distribution augments.
@@ IV Heuristics for modularity maximization
Maximizing the modularity function over all possible partitionings of the network is computationally hard, an NP-hard problem [42], hence the development of an array of heuristic algorithms [43, 9]. Here, we opt for the spectral method [7], the so-called leading eigenvector method. Initially designed for finding exactly two groups of nodes in the network in the most simple setting, spectral bipartitioning can easily be extended to an arbitrary number of communities by repeatedly dividing the communities into two until no modularity gain is to be achieved by so doing.
@@ IV.1 Spectral partitioning
Modularity matrix is, however, dense and, for example, the explicit matrix-vector multiplication $\mathbf{B}\mathbf{x}$ requires $O(|V|^{2})$ operations. Ref [7] shows how the matrix-vector multiplication $\mathbf{B}\mathbf{x}$ with a simpler definition of modularity can be rewritten in vector notation to allow for fast computation. That derivation is not directly applicable to our more complicated block-corrected null model, and we next show to derive a similar procedure for it. We can take advantage of the definition of $B$ and the fact that typically adjacency matrix $\mathbf{A}$ is sparse to write it as $\mathbf{B}\mathbf{x}=\mathbf{A}\mathbf{x}-\mathbf{P}\mathbf{x}$. Recalling edge probability in Eq. (2), we define
@@
Writing this in terms of element-wise multiplication operator $\odot$ gives
@@ IV.3 Final-Louvain fine-tuning
In addition to the split-Louvain fine-tuning, we implement another method of fine-tuning akin to phase one of the Louvain algorithm [9], that is applicable to a partition with an arbitrary number of communities. Unlike the split-Louvain fine-tuning nodes can be moved between communities multiple times, so long as the corresponding $\Delta Q$ values are positive. While in split-Louvain fine-tuning the best possible move among all nodes is selected in each iteration, in the multi-community final-Louvain fine-tuning we perform the most beneficial move for every node on an arbitrary order of nodes in each round and it is repeated until no modularity-gaining move remains for any of the nodes. Unlike the split-Louvain fine-tuning which is applied after every bipartition, we only perform the final-Louvain fine-tuning on the final output partition of the community detection algorithm.
@@ V.1 Field networks
We select eight networks ranging in size and scientific discipline as summarized in Table 1. Nodes in the networks are articles whose publication dates span from the beginning of the year 2000 to 2025, removing all whose in-degrees are less than 10. We temporally bin nodes by year, resulting in 24 time layers in total. Since the networks need not be directed acyclic, we allow for edges within time layers or against the arrow of time.
@@
For all networks in Table 1 the split-Louvain fine-tuning results in larger modularity values than the final-Louvain. One way to explain this is by the same phenomenon of error accumulation that, we posit, accounts for the fluctuations across community detection outcomes of a network. Split-Louvain fine-tuning is executed after every bipartition and thus diminishes the error that might compound over all subsequent iterative steps. In contrast, final-Louvain fine-tuning is performed only on the final outcome partition. Although applying both types of fine-tuning only marginally improves the results of split-Louvain, it typically adds a negligible amount of extra computation time compared to using split-Louvain alone.
@@ V.2 Subfield networks and metadata alignment evaluation
Biochem. – Clinical biochem. Environ. chem. – Analytical chem. Environ. chem. – Ocean eng. Environ. eng. – Pollution
@@
Pharma. – Social psychology Polymers – Biomaterial Aerospace eng. – Mechanics of material Physiology – Cardiology med.
@@
Unlike synthetic networks, there are generally no planted communities in real-world networks, nor the true generative process is known [47]. Absent a ground-truth partition, metadata, typically node attributes, is often treated as such. The assumption, nevertheless, that metadata is inherently linked with the community structure cannot be readily made [47, 48, 49]. Alongside the deep-seated reasons metadata does not always correlate with network structure, the process of generating metadata labels can be noisy and/or biased. In the OpenAlex database, for instance, citation relationship is a feature for the classifier that produces the node labels. With said caveats in mind, for randomly selected networks we evaluate whether partitions found by our model and the DAG [14] bear any resemblance to that of the metadata partitions.
@@
For each modularity, we compute a global $F_{1}$ score (based on the total number of true positives, false negatives and false positives) indicating how well they recover the two original subfields, to inspect how the $F_{1}$ scores of the two models compare for networks with different levels of separatedness, defined as the fraction of out-group edges to all edges. To illustrate this, in Fig. 6, we plot the probability that Block-corrected performs better than DAG for different maximum thresholds of separatedness. As the separatedness of the two subfields accentuates, the labeling advantage of block-corrected compared to DAG increases. This shows that the block-corrected modularity is more effective at recovering the partitions imposed by the omitted (unknown) subfield metadata from the network structure compared to DAG, especially in networks where innate communities imposed by subfields are more distinct.
Figure 6: Probability of the DAG model producing outcomes closer to the metadata compared to the block-corrected, as measured by $F_{1}$ scores, decreases with the minimum threshold on separatedness of the metadata communities. The $F_{1}$ scores are produced by performing spectral bipartitioning, followed by Split fine-tuning (see Sec. IV.2). Note that the separatedness on the horizontal axis is plotted cumulatively. Shaded area shows 95% confidence interval.

### v2

@@ I Introduction
Here, we take the opposite approach: for a given network data accompanied by node metadata, we discount the effect of known attributes on network structure in order to isolate the effect of unknown attributes. This amounts to asking whether there exists any community structure beyond the known attributes. Factoring out the known attributes has been previously accomplished in spatial networks, for example, by adapting the definition of modularity to discount the effects of spatial embedding [17]. In a mobile phone network where calls are predominantly placed among persons in close physical proximity, nullifying this physical proximity of callers came to reveal linguistically homogeneous communities, communities of callers speaking the same language. Similar approaches have been proposed for temporal networks by accounting for an ordering on the nodes, in either modularity-based [14] or Bayesian inference methods [29].
@@ III.1 Simple network model with hidden communities
where edge probability is proportional to the in-degree and out-degree of nodes, normalized by the total number of edges. We analytically compute modularity values of specific partitions with respect to both known and unknown structures to assess which partition is preferred by the block-corrected modularity and the directed modularity.
@@
The derived modularity expressions for $Q^{\mathrm{dir}}_{y}$, $Q^{\mathrm{bc}}_{y}$ and $Q^{\mathrm{dir}}_{x}$ have a similar form to the E-I index [42] which equates to the number of internal links of group members subtracted from the number of external links divided by their summation. Similar to the E-I index, modularity values $Q_{x}$ and $Q_{y}$ are measures of the dominance of in-group over out-group edges. As per Eq. (4) and Eq. (5) for a given binary attribute the more segregated its two membership groups are, the larger its corresponding modularity value is. The block-corrected modularity evidently rejects the known attribute community structure ($Q^{\mathrm{bc}}_{x}=0$) regardless of how well-separated and embedded its two comprising groups, i.e., the nodes with $x=0$ and those with $x=1$, are. In contrast, the directed null model’s ability to recover the unknown structure depends on the separatedness of the corresponding attribute’s groups. If the separation of the two groups under known attribute $x$ is more pronounced than those of the unknown attribute $y$, the modularity value $Q_{x}$ is larger than $Q_{y}$, in which case a modularity optimization partitioning algorithm will dismiss the community structure driven by the unknown attribute $y$ in favor of partitioning the network based on the known attribute $x$. The details of the derivations are outlined in Appendix A.
While it may seem plausible to approach this problem by partitioning the network into four communities, i.e., one community for each combination of $x$ and $y$, it is noteworthy that such a partition would only be preferred under certain combinations of parameters $p_{0}^{x}$, $p_{1}^{x}$, $p_{0}^{y}$ and $p_{1}^{y}$. In Appendix A we present a combination of parameters where a four-community partition would not be the solution, while the method presented in this paper provides a more general solution for nullifying any pattern of connectivity imposed by known attributes.
@@ III.2 Communities in temporal networks
A null model is characterized by its edge probability $P_{ij}$, which describes a generative network model for networks maximally similar to an observed network with no community structure present. Null models may nonetheless exhibit structures different from those existing in the observed network which can speak to their efficacy as a generative model. With citation networks in mind, we evaluate the performance of our proposed block null model against two previously proposed null models: the directed null model [41], which we already introduced in the previous subsection, and the DAG null model [14, 36]. The directed and the DAG null models are designed to operate on directed networks, with the latter specifically devised with citation networks in mind as they are often represented by directed acyclic graphs [^1] Real-world citation networks, however, are not always acyclic, requiring pre-processing measures to ensure the acyclic nature holds.. While the directed null model preserves only in- and out-degrees of nodes, the DAG null model additionally encodes two parameters $\mu_{t}$ and $\lambda_{t}$ per time layer $t$ describing the edge connectivity between layers:
@@ III.2.2 Temporal connectivity pattern in null models
To assess this, we define a single generalized $(bt_{\mathrm{max}})\times(bt_{\mathrm{max}})$ connectivity matrix $\mathbf{M}$ for each temporal network, where $M_{(bt_{i}+b_{i})(bt_{j}+b_{j})}$ is the fraction of existing edges (out of all possible edges) from community $b_{i}$ at time $t_{i}$ to community $b_{j}$ at time $t_{j}$. We then compare this to the connectivity matrix of the realizations of the same network passed through block-corrected, DAG, and directed null models.
@@
Figure 3: (a) Schematic of the skewed generative network model with two planted communities and a temporal structure where only nodes in adjacent blocks are connected. In addition, there are long links from the last block to the first. (b) Averaged connectivity matrices $\mathbf{M}$ of an ensemble of 100 network realizations of the block-corrected, DAG, and directed null models. The connectivity matrix of the original network, generated using the skewed model (Eq. (10)), is labeled “Original”. The skewed model network has $b=2$ planted communities, $t_{\mathrm{max}}=10$ temporal layers, $N=100$ nodes per temporal layer, mean in-group degree $\langle k^{\mathrm{ing}}\rangle=10$ and out-group degree $\langle k^{\mathrm{outg}}\rangle=8$. (c) The distribution of temporal edge lengths for all three network models: one with a skewed distribution of edge lengths (Eq. (10)), one with an exponential distribution (Eq. (11)), and lastly one with a power law edge length distribution (Eq. (12)). The three observed networks for skewed, exponential, and power law in panel (c) have $t_{\mathrm{max}}=50$ temporal blocks, i.e. time layers, with a fixed number of nodes ($N=100$) per temporal block. The number of planted communities is $b=2$ and each node has a mean in-group degree $\langle k^{\mathrm{ing}}\rangle=10$ and out-group degree $\langle k^{\mathrm{outg}}\rangle=8$. Edge length distributions for the drawn networks from the three null models are shown in their respective columns.
Averaged connectivity matrices of the realizations for the block-corrected, DAG, and directed null models are shown in the following three panels. The temporal connectivity patterns produced in our block-corrected model closely resemble those of the original skewed network, whereas the DAG model smooths the temporal connectivity pattern into an exponential decay. The directed model, on the other hand, assumes almost all community-time pairs are equally likely to be connected.
@@
Figure 3(c) shows that both the block-corrected and the DAG models recover the exponential edge length distribution observed in the original network. This is not the case, however, when the observed edge lengths follow a more heavily skewed distribution. The DAG model falls short of reproducing the distribution, producing an exponential edge length decay instead, while the block-corrected model generates an ensemble whose edge lengths closely follow the observed distribution. Similarly, under a power law distribution for the edge lengths, the DAG model generates networks with exponential edge length distributions instead, whereas the original power law distribution is recreated by the block-corrected model. In all three generative network models with exponential and skewed distributions, the directed model proves ineffective at capturing the nature of the distributions, producing almost constant temporal edge length distributions only affected by the finite number of temporal blocks $t_{\mathrm{max}}$.
@@ III.2.3 Susceptibility to erroneous partitioning
As we will describe in Sec. V, heavy-tail distributions of edge lengths is observed in certain real-world systems, including in citation networks. As we saw in Fig. 3(c), while our block null model can replicate the temporal connectivity pattern induced by a power law edge length distribution, the two other null models do not reproduce the underlying distribution. To understand the implications of this, we directly measure the ability of each modularity definition to recover the planted partitions in the space of all possible bipartitions of the network. To simplify this comparison, we assume the number of planted communities is known in all cases.
The quality of a partition can be quantified by the adjusted Rand score between that and the planted partition. Rand Score in its original form is a similarity measure for two partitionings, defined as
$R=\frac{\sum_{i<j}^{|V|}\eta_{ij}}{\binom{|V|}{2}}\,,$ (13)
where $\eta_{ij}$ is equal to one if both partitionings agree that $i$ and $j$ are in the same or different communities [43]. Adjusted Rand Score modifies this definition so that a score of zero represents performance no better than random labeling of nodes and one is a complete match between the candidate partitioning and the true one [44]:
$\frac{R-E[R_{\text{random}}]}{R_{\text{max}}-E[R_{\text{random}}]}\,,$ (14)
where $E[R_{\text{random}}]$ is the expected Rand Score of random labeling and $R_{\text{max}}$ the maximum possible value for Rand Score corresponding to a perfect match.
[…]
@@
The DAG modularity, however, alternatively assigns the majority of the nodes of groups of time layers to one of the two communities for smaller values of $\gamma$. This behavior, which is not reminiscent of the planted (original) communities, can already be seen at $\gamma=-2.0$ and seems to aggravate for more heavy-tailed distributions. This, again, corroborates the observations of a drop in entropy in Fig. 5(b). As hypothesized before, the directed modularity makes a vertical cut for all three values of $\gamma$ visualized in Fig. 5(c). The cut seems to become more gradual as the magnitude gets smaller, i.e., the heavy tail characteristic of the edge length distribution augments.
@@ IV Heuristics for modularity maximization
Maximizing the modularity function over all possible partitionings of the network is computationally hard, an NP-hard problem [45], hence the development of an array of heuristic algorithms [46, 9]. Here, we opt for the spectral method [7], the so-called leading eigenvector method. Initially designed for finding exactly two groups of nodes in the network in the simplest setting, spectral bipartitioning can easily be extended to an arbitrary number of communities by repeatedly dividing the communities into two until no modularity gain is to be achieved by so doing.
@@ IV.1 Spectral partitioning
Modularity matrix is, however, dense and, for example, the explicit matrix-vector multiplication $\mathbf{B}\mathbf{x}$ requires $O(|V|^{2})$ operations. Ref [7] shows how the matrix-vector multiplication $\mathbf{B}\mathbf{x}$ with a simpler definition of modularity can be rewritten in vector notation to allow for fast computation. This derivation is not directly applicable to our block-corrected modularity, and we next derive a similar procedure for it. We can take advantage of the definition of $B$ and the fact that typically adjacency matrix $\mathbf{A}$ is sparse to write it as $\mathbf{B}\mathbf{x}=\mathbf{A}\mathbf{x}-\mathbf{P}\mathbf{x}$. Recalling edge probability in Eq. (2), we define
@@
Writing this in terms of the element-wise multiplication operator $\odot$ gives
@@ IV.3 Final-Louvain fine-tuning
In addition to the split-Louvain fine-tuning, we implement another method of fine-tuning akin to phase one of the Louvain algorithm [9], that is applicable to a partition with an arbitrary number of communities. Unlike the split-Louvain fine-tuning nodes can be moved between communities multiple times, so long as the corresponding $\Delta Q$ values are positive. While in split-Louvain fine-tuning, the best possible move among all nodes is selected in each iteration, in the multi-community final-Louvain fine-tuning, we perform the most beneficial move for every node in an arbitrary order of nodes in each round and repeat it until no modularity-gaining move remains for any of the nodes. Unlike the split-Louvain fine-tuning which is applied after every bipartition, we only perform the final-Louvain fine-tuning on the final output partition of the community detection algorithm.
@@ V.1 Field networks
We select eight networks ranging in size and scientific discipline as summarized in Table 1. Nodes in the networks are articles whose publication dates span from the beginning of the year 2000 to 2025, removing all whose in-degrees are less than 10. We bin nodes in time by year, resulting in 24 time layers in total. Since the networks need not be directed acyclic, we allow for edges within time layers or against the arrow of time.
@@
For all networks in Table 1, the split-Louvain fine-tuning results in larger modularity values than the final-Louvain. One way to explain this is by the same phenomenon of error accumulation that, we posit, accounts for the fluctuations across community detection outcomes of a network. Split-Louvain fine-tuning is executed after every bipartition and thus diminishes the error that might compound over all subsequent iterative steps. In contrast, final-Louvain fine-tuning is performed only on the final outcome partition. Although applying both types of fine-tuning only marginally improves the results of split-Louvain, it typically adds a negligible amount of extra computation time compared to using split-Louvain alone.
@@ V.2 Subfield networks and metadata alignment evaluation
Biochem. – Clinical biochem. Environmental chem. – Analytical chem. Environmental chem. – Ocean eng. Environmental eng. – Pollution
@@
Pharmacology – Social psych. Polymers & plastics – Biomater. Aerosp. eng. – Mech. of material Physiology – Cardio. med.
@@
In addition to the information reported in Table 1, we include the modularity of the bipartition based on metadata labels (Metadata $Q$) and the modularity of the best bipartition of a randomized realization of the same network drawn from the block null model. The significant gap between $Q$ and $Q_{\text{randomized}}$ shows that the community structure present in the network is well beyond the detectability limit of this model. See Appendix E for more information on this issue.
@@
Unlike synthetic networks, there are generally no planted communities in real-world networks, nor the true generative process is known [50]. Absent a ground-truth partition, metadata, typically node attributes, is often treated as such. The assumption, nevertheless, that metadata is inherently linked with the community structure cannot be readily made [50, 51, 52]. Alongside the deep-seated reasons that metadata does not always correlate with network structure, the process of generating metadata labels can be noisy and/or biased. In the OpenAlex database, for instance, citation relationship is a feature for the classifier that produces the node labels. With said caveats in mind, for randomly selected networks we evaluate whether partitions found by our model and the DAG [14] bear any resemblance to that of the metadata partitions.
@@
For each modularity, we compute a global $F_{1}$ score (based on the total number of true positives, false negatives and false positives) indicating how well they recover the two original subfields, to inspect how the $F_{1}$ scores of the two models compare for networks with different levels of separatedness, defined as the fraction of out-group edges to all edges. To illustrate this, in Fig. 6, we plot the probability that Block-corrected performs better than DAG for different maximum thresholds of separatedness. As the separatedness of the two subfields accentuates, the labeling advantage of block-corrected compared to DAG increases. This shows that the block-corrected modularity is more effective at recovering the partitions imposed by the omitted (unknown) subfield metadata from the network structure compared to DAG, especially in networks where innate communities imposed by subfields are more distinct. Here the choice of $F_{1}$ score is justified since the problem at hand is a binary classification. However, in a more general setting, a more established metric for partition comparison such as the normalized mutual information is more fitting. We additionally repeated this analysis using normalized mutual information as the measure of performance with similar results, presented in Appendix D.
Figure 6: Probability of the DAG model producing outcomes closer to the metadata compared to the block-corrected, as measured by $F_{1}$ scores, decreases with the minimum threshold on separatedness of the metadata communities. The $F_{1}$ scores are produced by performing spectral bipartitioning, followed by split-Louvain fine-tuning (see Sec. IV.2). Note that the separatedness on the horizontal axis is plotted cumulatively. Shaded area shows 95% confidence interval.